Operation Summary for April 27, 2011 - April 29, 2011 |
|
Wednesday April 27 The day shift began with Operations monitoring Tevatron (TEV) store 8697 (96.74E30) and the Recycler stash (213.6E10), with the Antiproton Source (Pbar) stacking (20.0E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTFB T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations reported some short downtimes caused by Linac and Booster RF station trips. During the evening shift, Operations disabled beam to T-978 at 4:49 PM. Operations terminated store 8697 at 4:50 PM. RF personnel worked on a TEV RF station (TRF7) in between stores. Operations established HEP store 8698 at 6:04 PM with an average initial luminosity (AIL) of 385.5E30. TRF4 tripped off towards the end of shift causing a loss of 5E30 in luminosity. Operations reported some short downtime was caused by Linac and Booster RF station trips.Thursday April 28 The midnight shift began with Operations monitoring store 8698 (129.27E30), the stash (174.4E10), and the Pbar stack (12.3E10), and with NuMI and MiniBooNE waiting for Linac to recover. During the midnight shift, Operations reported that the I- Source output started drifting downward; operators tuned on the Source for most of the shift. Operations established beam to T-978 at 4:06 AM. Operations reported several Booster RF trips caused some minor downtime. During the day shift, Operations terminated store 8698 at 10:38 AM. RF personnel changed out the TRF4 modulator in between stores. TRF4 then developed a switch problem that held off shot setup for another hour. Operations established store 8700 at 1:55 PM with an AIL of 405.25E30. This store set a new TEV luminosity record. Operations reported that Linac and Booster RF work held off beam for about an hour after the shot. Proton Source beam resumed at 3:14 PM. During the evening shift, Operations disabled beam to T-978 at 6 PM. Operations monitored New Muon Lab coupler conditioning for cavity #4. Friday April 29 The midnight shift began with Operations monitoring store 8700 (119.24E30) and the stash (235.0E10), with Pbar stacking (18.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 at 4:15 AM. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 04/29/2011 email Fermilab |
| Security, Privacy, Legal | |